# Dance Like a Chicken: Low-Rank Stylization for Human Motion Diffusion

### v2

@@ 1 Introduction
Equipped with this insight, a simple solution, of employing Low-Rank Adaptation (LoRA) [26], becomes clear; LoRA augments the weights of the prior in a low-ranked manner, which can be intuitively thought of as low frequency deformations to the target distribution manifold. Using LoRA, LoRA-MDM learns to embed a style within the output domain of the model, which can then be used to synthesize stylized motions in the context of different textual prompts. As we demonstrate, this offers two advantages: First, low-ranked augmentation diminishes forgetting, which in turn increases generation quality for stylized actions not seen during training. For example, directly fine-tuning a generated walking motion to walk like a chicken has little (or negative) influence on how the prior will generate kicks. Instead, our low-ranked tuning and structure preservation loss encourage the model to also kick like a chicken, since the changes affect the entire manifold. Second, because the generated stylized motions are in-distribution for the model, advanced operations such as stylized motion editing and style blending are supported, in sharp contrast to traditional motion style-transfer method.
@@ 2.1 Human Motion Synthesis
In recent year, human motion generation has experienced significant progress with the advent of Transformers [66] and deep generative models [24, 15, 32]. Motions can be now generated from a variety of input modalities, including action categories [42, 17, 68, 38], music signals [2, 55, 34, 64], and textual descriptions [61, 18, 59, 72, 20, 19, 53, 6]. Early works achieved plausible 3D motion generation by leveraging variational autoencoders (VAEs) equipped with temporal models such as RNNs [18, 17, 34] and Transformers [42, 43]. However, these models often struggled to learn complex actions and handle fine-grained conditionality. To address this, a subsequent line of research modeled human motion generation in a discrete space. These approaches utilized VQ-VAE [65] to map human motions into discrete motion tokens, which were then generated from various modalities using language models like GPTs [70, 19, 55, 38] or generative masked Transformers [20, 44]. Diffusion models have also emerged as a powerful tool for motion synthesis due to their exceptional expressiveness. MDM [61] and MotionDiffuse [72] applied diffusion models directly in the raw motion feature space, while techniques like latent diffusion [6] and latent consistency models [9] have been adopted to enhance the efficiency of diffusion-based human motion synthesis.
@@ 3.2 LoRA-MDM
The method’s overview is illustrated in Figure 2. Our goal is to generate high-quality human motions in a specified style while preserving the diversity and generalization ability of a pre-trained motion diffusion model (MDM). Condition motion generation on style-specific data often leads to a degradation in motion quality and generalization due to the limited diversity of style datasets. On the other hand, fine-tuning the whole model using the style data, which includes a handful of motions per style, may lead to an overfit and the forgetting of underrepresented actions. Inspired by personalization literature [13, 50], we propose a lightweight adaptation method based on Low-Rank Adaptation (LoRA), which enables the model to internalize new styles while maintaining its learned motion manifold.
@@ 5.2 Style mixing
By embedding multiple style tokens within the same LoRA module, we enable style mixing —where conditioning on multiple tokens blends their stylistic attributes. Although each style is learned independently, combining tokens within a prompt results in a coherent fusion of styles. For example, a prompt like ”A person is […] in $\langle s_{1}\rangle$ style and in $\langle s_{2}\rangle$ style” generates a motion that seamlessly integrates both styles. Notably, not all style pairs can be mixed. For instance, hands above head and arms behind the back are physically conflicting and cannot be combined.
@@ 6 Discussion
In this paper we introduced LoRA-MDM, a method for motion stylization. We demonstrate that in the motion domain, similarly to images, a new concept can be generalized to unseen context through manifold preservation of a pre-trained prior. In essence, we showed that the manifold shift induced by low-rank adaptation outperforms in terms of quality and editability a mechanism that pushes the generation itself, as opposed to the model’s capabilities, towards a specific style. Furthermore, we hope this first demonstration of a LoRA employed for motion can guide when and how to use low-rank adaptation in this domain.
@@ C.1 Metrics
Top-$k$ accuracy of a trained style classifier, measuring how well the generated motions align with the intended styles. Since SMooDi employs the same style classifier for both classifier-based guidance and SRA evaluation, we used a separate classifier to ensure a fair comparison—see the details above.

### v3

@@ 1 Introduction
This insight motivates the use of Low-Rank Adaptation (LoRA) [26]. LoRA augments the weights of the prior in a low-ranked manner, which can be intuitively thought of as low frequency deformations to the target distribution manifold. Using LoRA, LoRA-MDM learns to embed a style within the output domain of the model, which can then be used to synthesize stylized motions in the context of different textual prompts. As we demonstrate, this offers two advantages: First, low-ranked augmentation diminishes forgetting, which in turn increases generation quality for stylized actions not seen during training. For example, directly fine-tuning a generated walking motion to walk like a chicken has little (or negative) influence on how the prior will generate kicks. Instead, our low-ranked tuning and structure preservation loss encourage the model to also kick like a chicken, since the changes affect the entire manifold. Second, because the generated stylized motions are in-distribution for the model, advanced operations such as stylized motion editing and style blending are supported, in sharp contrast to traditional motion style-transfer method.
@@ 2.1 Human Motion Synthesis
In recent year, human motion generation has experienced significant progress with the advent of Transformers [66] and deep generative models [24, 15, 32]. Motions can now be generated from a variety of input modalities, including action categories [42, 17, 68, 38], music signals [2, 55, 34, 64], and textual descriptions [61, 18, 59, 72, 20, 19, 53, 6]. Early works achieved plausible 3D motion generation by leveraging variational autoencoders (VAEs) equipped with temporal models such as RNNs [18, 17, 34] and Transformers [42, 43]. However, these models often struggled to learn complex actions and handle fine-grained conditionality. To address this, a subsequent line of research modeled human motion generation in a discrete space. These approaches utilized VQ-VAE [65] to map human motions into discrete motion tokens, which were then generated from various modalities using language models like GPTs [70, 19, 55, 38] or generative masked Transformers [20, 44]. Diffusion models have also emerged as a powerful tool for motion synthesis due to their exceptional expressiveness. MDM [61] and MotionDiffuse [72] applied diffusion models directly in the raw motion feature space, while techniques like latent diffusion [6] and latent consistency models [9] have been adopted to enhance the efficiency of diffusion-based human motion synthesis.
@@ 3.2 LoRA-MDM
The method’s overview is illustrated in Figure 2. Our goal is to generate high-quality human motions in a specified style while preserving the diversity and generalization ability of a pre-trained motion diffusion model (MDM). Conditioning motion generation on style-specific data often leads to a degradation in motion quality and generalization due to the limited diversity of style datasets. On the other hand, fine-tuning the whole model using the style data, which includes a handful of motions per style, may lead to an overfit and the forgetting of underrepresented actions. Inspired by personalization literature [13, 50], we propose a lightweight adaptation method based on Low-Rank Adaptation (LoRA), which enables the model to internalize new styles while maintaining its learned motion manifold.
@@ 5.2 Style mixing
By embedding multiple style tokens within the same LoRA module, we enable style mixing - where conditioning on multiple tokens blends their stylistic attributes. Although each style is learned independently, combining tokens within a prompt results in a coherent fusion of styles. For example, a prompt like ”A person is […] in $\langle s_{1}\rangle$ style and in $\langle s_{2}\rangle$ style” generates a motion that seamlessly integrates both styles. Notably, not all style pairs can be mixed. For instance, hands above head and arms behind the back are physically conflicting and cannot be combined.
@@ 6 Discussion
In this paper we introduced LoRA-MDM, a method for motion stylization. We demonstrate that in the motion domain, similarly to images, a new concept can be generalized to unseen context through manifold preservation of a pre-trained prior. In essence, we showed that the manifold shift induced by low-rank adaptation outperforms in terms of quality and editability. Furthermore, we hope this first demonstration of a LoRA employed for motion can guide when and how to use low-rank adaptation in this domain.
@@ C.1 Metrics
Top-$k$ accuracy of a trained style classifier, measuring how well the generated motions align with the intended styles. Since SMooDi employs the same style classifier for both classifier-based guidance and SRA evaluation, we used a separate classifier to ensure a fair comparison-see the details above.